40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11197     Entry time: Fri Apr 3 05:17:36 2015
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI to 1f twelve times 

I have 12 tick marks for times I got all the way to 1f for all 4 degrees of freedom in the PRFPMI.  The CARM / DARM transitions now succeed more than they fail, which is nice. 

At Q's suggestion, I am turning off all the violin filters in the MC2 path during the CARM transition.  This also means that I don't need any of the notches that Den and I put into the CARM_A and CARM_B filter banks last week, which were right at the edges of the violin notches.  Anyhow, this seems to make the transition much more likely to succeed.  I don't ever use the CM_SLOW FM10 "crossover helper" that Q had to use last night.  The violin filters are turned back on after the CARM transition is complete.  We don't ever need those other notches.

I checked the REFL165 vs. REFL55 transfer functions for PRCL and MICH, and they are mostly flat.  REFL55 seems like it'll give us extra phase for some reason.

I tried setting offsets for PRMI, but they seem to be strongly dependent on arm alignment.  I ended up being pretty confused, and since all the REFL signals are pretty close to zero (when CARM/DARM on RF, PRMI on 3f), I have given up on that avenue for tonight. 

I think many of my locklosses tonight (lost from the all 1f state) have been fast things, faster than the ADCs can handle.  On the lockloss plots that I've looked at, the FSS PC drive is railed at 10V about 200msec before I lose lock.  So, something (presumably in the fast CARM path) is making the MC/FSS loop unhappy.  I have plugged in the Agilent to the Out2 of the CM board, so that it looks at REFL11.  Unfortunately, this is after the input gain slider, so we don't see much until we're locked, but that seems fine.  A video camera is pointed at the screen, so that I get real time spectra.  It's hard to watch the TV at the same time as everything else, so I haven't witnessed the moment of lockloss in the fast spectrum yet.  Be careful when walking down the Yarm.  The tripod is partly in the walkway.

Q, I took a few TFs of the total CARM loop, although none of them are particularly good below a few kHz.  I can't push hard enough to get coherence, without blowing the lock.  TF data is in /users/jenne/PRFPMI/CM_TFs/CM_TFs_2Apr2015/ . 

I was worried for a while that, after I transition PRMI to 1f, I hear lots of low frequency rumbling.  However, watching the spectra (relative to references taken with CARM and DARM on RF, but PRMI on 3f), the low frequency error and control signals are staying the same for all 4 DoFs, but the high frequency for PRCL and MICH goes down significantly, so it's probably just that the low frequency stuff sounds more obvious, since it's not drowning in high frequency fuzz. 

 

ELOG V3.1.3-